let the function f definedbe defined as: f(x)= 4a if x</=-2, 3x^2 if -2<x</=1, x+b if x>1. Determine the values of the consonants a,b, so that f is continuous at x=-2 & x=1
1
Expert's answer
2021-07-19T05:50:04-0400

At the point x=-2: "4a=3*(-2)^2 \to a=3."

At the point x=1: "3*1^2=1+b\to b=2."